OpenStudy (hapahearthero):

I'm not sure what this question is asking.. Says find and simplify the expression f(x)=x^2-7 then below it says: f(2+h)-f(2)=? (calculus class)

OpenStudy (peachpi):

find f(2 + h) by substituting 2 + h for x in f(x). Find f(2) by plugging in 2 for x. subtract

OpenStudy (peachpi):

f(2 + h) = (2 + h)² - 7 f(2) = 2² - 7 f(2 + h) - f(2) = (2 + h)² - 7 - (2² - 7)

OpenStudy (hapahearthero):

Thank you I got it!!!